# Quantitative propagation of chaos for 2D stochastic vortex model on the whole space under moderate interactions

### v1

@@ Abstract
We derive the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in the sense of relative entropy. The main novelty is the application of the Donsker–Varadhan inequality in the context of moderately interacting particles to handle the nonlinearity, as well as the use of localization techniques combined with the probabilistic data setting, to derive estimates for the quadratic variation terms. Moreover, we prove the existence of a suitable solution to the aforementioned model.
@@ 1 Introduction
In order to derive (5), we apply for the first time in the context of moderately interacting particles, the Donsker–Varadhan inequality together with the dissipation provided by the Fisher information, to handle the nonlinearity. In addition, we employ localization techniques combined with the probabilistic data setting, to derive estimates for the quadratic variation terms in the whole space, as the estimate established in [39], under periodic boundary conditions, does not apply in the present framework. Moreover, we prove the existence of a suitable solution to the equation (1).
@@ 1.3 Statement of the main results
$\lim_{N\to\infty}N^{\theta}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.$
@@
$\theta=\min\left(1-\beta(2+\frac{2}{d}+2\alpha);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
with $\delta>0$, such that $\theta>0$, $d\geq 1$ and $\beta\in\left(0,\frac{1}{2\big[1+\frac{1}{d}\big]}\right)\cap\left(0,\frac{1}{\big[2+\frac{2}{d}+2\alpha\big]}\right)$.
Then, there exists a time $T\in(0,T_{1})$, such that
$\lim_{N\to\infty}N^{\theta}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})=0,\,\,\mathbb{P}-a.s.$
@@
$\lim_{N\to\infty}N^{\theta-\delta}\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}=0,\,\,\,\mathbb{P}-a.s.$
@@ 2.2 Estimates for $II_{t}+IV_{t}$: quadratic variation terms
$II_{t\wedge\tau^{N}}+IV_{t\wedge\tau^{N}}$ $=\frac{1}{2N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{1}{\rho_{s}^{N}(x)}|\nabla V^{N}(x-X_{s}^{i,N})|^{2}\,dx\,ds$
$\overset{}{\leq}\frac{N^{\beta}N^{\frac{2\beta}{d}}}{2N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{1}{\rho_{s}^{N}(x)}|V^{N}(x-X_{s}^{i,N})|^{2}\,ds\,dx$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\frac{\frac{1}{N}\sum_{i=1}^{N}V^{N}(x-X_{s}^{i,N})}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
$\overset{\alpha>d/2}{\leq}C_{\alpha,T}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}.$ (36)
@@ 2.3 Estimates for $I_{t}$: the nonlinear terms
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{\eta>>1}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\eta\ln{(1)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s}),$ (43)
@@
$I_{t}^{1}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+\epsilon\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds.$ (45)
@@
$\overset{}{\leq}2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx,$ (48)
@@
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{\eta>>1}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\eta\ln{(1)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s}).$
@@
$\leq 1/2\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}},$
@@
$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\,dx\,ds$
@@
$\overset{\text{Holder}}{\leq}(\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds.$ (49)
@@
$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds.$ (50)
@@ 2.4 Application of Gronwall’s Lemma
$\overset{}{+}(2\epsilon+\sqrt{2}/2)\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
@@
$+C_{\alpha,T}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}$
$+M_{t}^{N}.$ (51)
With the same computations as in [39], for all $\delta\in(0,\theta_{1})$ there exists a random variable $A_{0}$, with finite moments such that
$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{1}+\delta},$
which allows us to express (51), for $\epsilon<<1$, as
$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+C_{\alpha,T}N^{-\theta_{2}}+A_{0}N^{-\theta_{1}+\delta},$ (52)
with $\theta_{2}\doteq 1-\beta(2+\frac{2}{d}+2\alpha)$.
From Gronwall’s Lemma in (52), recalling (26), (30) and (29), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}\right)$ (53)
@@
$\theta\doteq\min\left(1-\beta(2+\frac{2}{d}+2\alpha);\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta\right).$
@@ 2.5 End of the proof: Removing the stopping time
Now we will derive an analogous estimate for $L_{3}^{N}$. First, we observe that, if $\sigma=0$ we are done, thus suppose $\sigma\neq 0$. Since $\sigma$ is bounded, the function defined by $t\mapsto\int_{0}^{t}\sigma^{2}_{s}\,ds$ is continuous and increasing. Hence, recalling (18), by Theorem 4.6, in Chapter 3 of [25], for all $j\in\{1,...,d\}$, we get
$\sup_{t\leq T}|X_{t}^{j}|=\sup_{t\leq T}|\widetilde{B}^{j}_{\langle X\rangle_{t}}|\leq\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j}_{t}|,$ (62)
where $(\widetilde{B}_{t})_{t\geq 0}$ is the time-change, which in particular, is a one dimensional standard Brownian motion. In addition, since
@@
and by Borel-Cantelli Lemma, we arrive at
@@
$\lim_{N\to\infty}N^{\theta}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})=0,$

### v2

# arXiv:2602.22038v2[math.AP] 27 Feb 2026

@@ Abstract
We derive quantitative estimates for the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in the terms of relative entropy. The main novelties lie in the application of the Donsker–Varadhan inequality within the framework of moderately interacting particle systems to handle the nonlinear terms, as well as in the introduction of localization techniques combined with a probabilistic data setting, in order to derive estimates for the quadratic variation terms. Moreover, we prove the existence of a suitable solution to the aforementioned model.
@@ 1 Introduction
In the present work, we apply for the first time in the context of moderately interacting particles, the Donsker–Varadhan inequality together with the dissipation provided by the Fisher information, to handle the nonlinearity. In addition, we employ localization techniques combined with the probabilistic data setting, to derive estimates for the quadratic variation terms in the whole space, as the estimate established in [39], under periodic boundary conditions, does not apply in the present framework. Moreover, we prove the existence of a suitable solution to the equation (1).
@@ 5. $(\mathbf{A}^{X_{0}})$
Remark 3.
Assumption $(\mathbf{A}^{\rho_{0}})$ was formulated in [7, 48, 49], where, relying on Li–Yau–type gradient bounds and Hamilton–type heat kernel estimates, the authors use it to control the solution and its derivatives up to second order. It allows the relative entropy approach of [23] to be extended from the periodic framework to the whole space.
@@ 1.3 Statement of the main results
$\mathcal{H}(\rho_{0}^{N}|\rho_{0})\leq C,\,\,\,\mathbb{P}-a.s.$
@@
$\theta=\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
with $\delta>0$, such that $\theta>0$, $d\geq 1$ and $\beta\in\left(0,\frac{1}{2\big[1+\frac{1}{d}\big]}\right)\cap\left(0,\frac{1}{\big[1+\frac{2}{d}+2\alpha\big]}\right)$.
Then, there exist a time $T\in(0,T_{1})$ and a constant $C>0$, such that, almost surely for $N>>1$,
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (19)
@@
$\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (20)
@@ 2.2 Estimates for $II_{t}+IV_{t}$: quadratic variation terms
[…]
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\frac{\frac{1}{N}\sum_{i=1}^{N}V^{N}(x-X_{s}^{i,N})}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
$\overset{\alpha>d/2}{\leq}C_{\alpha,T}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}.$ (38)
@@ 2.3 Estimates for $I_{t}$: the nonlinear terms
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}2\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta},$ (45)
@@
$I_{t}^{1}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+\epsilon\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (47)
@@
$\overset{}{\leq}2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx,$ (50)
@@
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}2\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}.$
@@
$\leq 1/2\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}},$
@@
$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds$
@@
$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds$
@@
$+\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+TC_{\eta}$
$\overset{\text{Holder}}{\leq}(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds+TC_{\eta}.$ (51)
@@
$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(3\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (52)
@@ 2.4 Application of Gronwall’s Lemma
In this subsection, we conclude our estimates, by applying Gronwall’s inequality.
Indeed, with the same computations as in [39], for all $\delta\in(0,\theta_{1})$ there exists a random variable $A_{0}$, finite almost surely, such that
$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{1}+\delta},$
with $\theta_{1}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
@@
$\overset{}{+}(3\epsilon+\sqrt{2}/2)\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
@@
$+C_{\alpha,T}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}$
$+A_{0}N^{-\theta_{1}+\delta}$
$+TC_{\eta}$ (53)
and then for $\eta>>1$ and $\epsilon<<1$, we deduce
$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+N^{-\theta_{2}}+A_{0}N^{-\theta_{1}+\delta}+C,$ (54)
with $\theta_{1}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$ and $\theta_{2}\doteq 1-\beta(1+\frac{2}{d}+2\alpha)$.
Finally, from Gronwall’s Lemma in (54), recalling (28), (31) and (32), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (55)
@@
$\theta\doteq\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta\right).$
@@ 2.5 End of the proof: Removing the stopping time
Now we will derive an analogous estimate for $L_{3}^{N}$. First, we observe that, if $\sigma=0$ we are done, thus suppose $\sigma\neq 0$. Since $\sigma$ is bounded, for $j\in\{1,...,d\}$, the function defined by $t\mapsto\langle X^{j}\rangle_{t}$ is continuous and increasing. Hence, recalling (18), by Theorem 4.6, in Chapter 3 of [25], for all $j\in\{1,...,d\}$, we get
$\sup_{t\leq T}|X_{t}^{j}|=\sup_{t\leq T}\left|\widetilde{B}^{j}_{\langle X^{j}\rangle_{t}}\right|\leq\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j}_{t}|,$ (64)
where $(\widetilde{B}^{j}_{t})_{t\geq 0}$ is the time-change process, which in particular, is a one dimensional standard Brownian motion. In addition, since
@@
and then by Borel-Cantelli Lemma, we arrive at
@@
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (70)